OpenStudy (anonymous):

1.) Simplify completely: the quantity 4 times x to the third power plus 5 times x to the 2nd power plus 3 times x all over x 2.) Simplify completely: the quantity 5 times x to the third power plus 10 times x to the 2nd power plus 15 times x all over 5 times x 3.) Simplify completely: the quantity 10 times x to the 6th power times y to the third power plus 20 times x to the third power times y to the 2nd power all over 5 times x to the third power times y Cqaan someone please help me??

jhonyy9 (jhonyy9):

hi

OpenStudy (anonymous):

hi

jhonyy9 (jhonyy9):

so i like help you but you need cooperate ,ok ?

OpenStudy (anonymous):

ok

jhonyy9 (jhonyy9):

1. 4x^3 +5x^2 +3x ---------------- = x so the first wann being this ,yes ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

so for you can simplifie anything you need factorizing out in numerator the common factor so what will be ?

jhonyy9 (jhonyy9):

in numerator there are 3 terms so what is common ?

OpenStudy (anonymous):

4x^3+5x^2?

jhonyy9 (jhonyy9):

no 4x^3 +5x^2 +3x what is common in this 3 ?

OpenStudy (anonymous):

i don't understand

jhonyy9 (jhonyy9):

4x^2*x +5x*x +3x so what is common here ?

OpenStudy (anonymous):

the varibles?

jhonyy9 (jhonyy9):

the x you see in every 3 terms yes ? so can you factorizing out ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

ok than what will get ?

OpenStudy (anonymous):

12x^2

jhonyy9 (jhonyy9):

no 4x^3 +5x^2 +3x = x(4x^2 +5x +3) ok ?

OpenStudy (anonymous):

what about the exponents?

jhonyy9 (jhonyy9):

x(4x^2 +5x +3) ---------------- = x so now you can simplifie the x and will get ..... what ?

jhonyy9 (jhonyy9):

what is understandably for you ?

OpenStudy (anonymous):

nothing really

jhonyy9 (jhonyy9):

why ? you dont know how you calculi with exponents ?

OpenStudy (anonymous):

yeah. 3-2=1

jhonyy9 (jhonyy9):

a^2 +a^2 =2a^2 a*a=a^2 a^2 *a =a^3

OpenStudy (anonymous):

I still don't understand. How do you even get a answer with a variable at the bottom.

jhonyy9 (jhonyy9):

so because in these 3 terms of numerator there is x so you can factorizing out it and will get x(4x^2 +5x +3) and getting this you can simplifie the x from numerator by x from denominator ok ?

OpenStudy (anonymous):

okay

jhonyy9 (jhonyy9):

sure ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

so than try simplifie the second exercise

jhonyy9 (jhonyy9):

i am here to check it , ok ?

OpenStudy (anonymous):

okay

jhonyy9 (jhonyy9):

???

OpenStudy (anonymous):

5x^3+10x^2=15x^1? 15x^1+15/5x?

OpenStudy (anonymous):

6?

jhonyy9 (jhonyy9):

no first try writing what will be in numerator

OpenStudy (anonymous):

15+15=30 30/5=6

OpenStudy (anonymous):

let me try again

jhonyy9 (jhonyy9):

please you write here the numerator

jhonyy9 (jhonyy9):

5x^3+10x^2+15x so this is the numerator

jhonyy9 (jhonyy9):

and 5x is the denominator right ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

so what is the common factor in numerator what you can factorizing out ?

OpenStudy (anonymous):

5x^3+10x^2

jhonyy9 (jhonyy9):

5x*x^2 +5x*2x +3*5x what is common in every 3 terms ?

OpenStudy (anonymous):

i dont understand

jhonyy9 (jhonyy9):

5x yes ? than you can factorizing out it and will get what ?

jhonyy9 (jhonyy9):

5x*x^2 = 5x^3

jhonyy9 (jhonyy9):

5x*2x=10x^2

OpenStudy (anonymous):

wait what question are you doing?

jhonyy9 (jhonyy9):

so do you understand it now ?

OpenStudy (anonymous):

what question are you doing?

jhonyy9 (jhonyy9):

sorry but i dont understand you

OpenStudy (anonymous):

are you explaining 2 or 3 to me?

jhonyy9 (jhonyy9):

2

OpenStudy (anonymous):

okay

jhonyy9 (jhonyy9):

so than ?

jhonyy9 (jhonyy9):

sorry but i need to go

OpenStudy (anonymous):

okay thank you for trying to help me

jhonyy9 (jhonyy9):

but you will can solve it sure ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

ok good luck bye

OpenStudy (anonymous):

thank you. bye!!